If vectors 2i+4j-7k and 2i+6j+qk are perpendiculars then value of q is?


1
Expert's answer
2021-08-25T17:24:03-0400

If vectors are perpendicular, their scalar producti is equal to zero. Thus, obtain the following equation:


"2\cdot 2 + 4\cdot 6 -7\cdot q = 0\\
q = \dfrac{28}{7} = 4"

Answer. 4.